OpenStudy (anonymous):

Ryan flew from Wiley Post to Ponca City and back. Ryan maintained an average rate of 450 mph going to Ponca City and an average rate of 400 mph returning to Wiley Post. If the actual flying time for the roundtrip was 1 hour, about how far is it from Wiley Post to Ponca City? Round to nearest mile

OpenStudy (matt101):

Let T be the time it took to go from WP to PC and t be the time it took to go back. Also remember the equation v = d/t. Since the distance is the same regardless of which way you go, you know 450T = 400t. Since the round trip was 1 hour, T + t = 1 You have two equations and two unknowns - solve :)

OpenStudy (anonymous):

this is so difficult

OpenStudy (matt101):

do you understand how I got the two equations?

OpenStudy (anonymous):

no idea

OpenStudy (matt101):

I'll use capital letters (D, V, T) to represent the first trip and lower case letters (d, v, t) to represent the second. The second equation is a little easier to come up with. The total time it takes for the round trip is 1 hour, so the time it takes to go one way (T) plus the time it takes to go the other way (t) must equal 1: T + t = 1. For the first equation, you know the equation velocity = distance/time (v = d/t). The distance between WP and PC is the same, regardless of which way you go. From v = d/t, you can find d by multiplying v and t. Since D = d, you know that VT = vt too. They give you V = 450 and v = 400 in the question, so 450T = 400t. Once you find T and t, you can use either one and the corresponding velocity to find the distance!

OpenStudy (anonymous):

so the answer is 212 miles

OpenStudy (matt101):

Yup :) Did you get how to solve it?